OpenStudy (nuccioreggie):

Solve for x: one over five (5x + 12) = 18 seventy-eight over five one hundred-two over five 6 30

OpenStudy (anonymous):

The equation is \(\dfrac{1}{5}\)(5x + 12) = 18, correct?

OpenStudy (anonymous):

@nuccioreggie

OpenStudy (nuccioreggie):

I got C

jagr2713 (jagr2713):

@nuccioreggie try again, you still there

OpenStudy (anonymous):

\[\frac{ 1 }{ 5 }\left( 5x + 12 \right) = 18\] so move this 5 to the right hand side

OpenStudy (anonymous):

i.e 5x + 12 = 18 times 5

OpenStudy (anonymous):

5x + 12 = 90

OpenStudy (anonymous):

5x = 90 -12

OpenStudy (anonymous):

5x = 78

jagr2713 (jagr2713):

ok first we multiply both sides by a constant to simplify the equation. multiply both sides of 5x+12/5= 18 by 5 \[\frac{ 5(5x+12) }{ 5 }= 5 x 18\]

OpenStudy (anonymous):

x = 78/5

jagr2713 (jagr2713):

@KenichiIvy please dont do the work for them. Thank you :D

jagr2713 (jagr2713):

what do you think we do next @nuccioreggie

OpenStudy (anonymous):

I'll take care NEXT time :)